1) Reeves Co. filed suit against Higgins, Inc., seeking damages for copyright violations. Higgins' legal counsel believes it is probable that Higgins will settle the lawsuit for an estimated amount in the range of $160,000 to $260,000, with all amounts in the range considered equally likely. How should Higgins report this litigation?

2)At the beginning of 2021, Angel Corporation began offering a 1-year warranty on its products. The warranty program was expected to cost Angel 4% of net sales. Net sales made under warranty in 2021 were $260 million. Five percent of the units sold were returned in 2021 and repaired or replaced at a cost of $9.0 million. The amount of warranty expense in Angel's 2021 income statement is:

3)Strikers, Inc. sells soccer goals to customers over the Internet. History has shown that 5% of Strikers' goals will need repair under the warranty program. For the year, Strikers has sold 4,100 goals and 47 have been repaired.

If the estimated cost to repair a goal is $120, what would be the Warranty Liability at the end of the year?

4) Talks-A-Lot, Inc. sells cell phones to customers and expects that 10% of phones sold will be returned for repair under its warranty program. The average repair cost is $75 per phone. For 2021, Talks-A-Lot has sold 590 cell phones and has repaired 11 of them as of December 31, 2021. What amount of warranty liability should be reported at December 31, 2021? (Do not round intermediate calculations.)
Answer #1

1.

Answer:

Report as a liability for $160,000 with disclosure of range

Explanation

Here we report 160,000 which is the minimum amount in the range as liability because all amounts in the range are treated equally likely. So no amount comes more likely than others as it is equally likely.

2.

Answer:

Amount of warranty expense in Angel's 2021 income statement = 10,400,000

Calculation:

Amount of warranty expense in Angel's 2021 income statement

Warranty Expense = Net sales * Cost % of net sales

So,

Warranty Expense = 260,000,000 * 4% = 10,400,000

3.

Answer:

Warranty Liability for the year = $ 18,960

Calculation:

$4,100 goals x 5% cost to repair warranty expense = 205

Warranty Expense = 205 per goals x $120 per goal = $24,600

Repair * Cost to repair = 47 * 120 = 5,640

We need to deduct this amount from Warranty Expense.

So, Warranty Liability = 24,600 - 5,640 = 18,960

4.

Answer:

Warranty liability at December 31, 2021 = 3,600

Calculation:

Estimated Loss = Repair units * Repair Cost

So, repair units = 590 units sold x 10% to repair = 59 units

Estimated Loss = 59 units *  $75 per unit to repair = 4,425

Then calculate Amount of actually repaired:
Amount actually repaired = 11 units repaired * $75 per unit = 825

So to find Ending balance of Warranty Liability:
Ending balance of Warranty Liability = Estimated Loss - Amount actually repaired

= 4,425 - 825 = 3,600